### v1

# Spatial structure facilitates evolutionary rescue by cost-free drug resistance

Abstract
Bacterial populations often have complex spatial structures, which can impact their evolution. Here, we study how spatial structure affects the evolution of antibiotic resistance in a bacterial population. We consider a minimal model of spatially structured populations where all demes (i.e., subpopulations) are identical and connected to each other by identical migration rates. We show that spatial structure can facilitate the survival of a bacterial population to antibiotic treatment, starting from a sensitive inoculum. Indeed, the bacterial population can be rescued if antibiotic resistant mutants appear and are present when drug is added, and spatial structure can impact the fate of these mutants and the probability that they are present. Specifically, if the mutation that provides resistance is neutral or effectively neutral, its probability of fixation is increased in smaller populations. This promotes local fixation of resistant mutants in the structured population, which facilitates evolutionary rescue by cost-free drug resistance. Once the population is rescued by resistance, migrations allow resistant mutants to spread in all demes. Our main results extend to the case where there are resistant mutants in the inoculum, and to more complex spatial structures. They also extend to resistant mutants that carry a fitness cost, although the timescales involved are longer.
Author Summary
Antibiotic resistance is a major challenge, since bacteria tend to adapt to the drugs they are subjected to. Understanding what conditions facilitate or hinder the appearance and spread of resistance in a bacterial population is thus of strong interest. Most natural microbial populations have complex spatial structures. This includes host-associated microbiota, such as the gut microbiota. Here, we show that spatial structure can facilitate the survival of a bacterial population to antibiotic treatment, by promoting the presence of resistant bacteria. Indeed, neutral mutants giving resistance can take over small populations more easily than large ones, thanks to the increased importance of fluctuations in small populations. Resistant mutants can then spread to the whole structured population. Thus, population spatial structure can be a source of antibiotic treatment failure. This effect of spatial structure is generic and does not require environment heterogeneity.
Introduction
Antibiotic resistance is a crucial challenge in public health [who_antibiotics2, eu_ab]. Resistant bacteria emerge and spread through Darwinian evolution, driven by random mutations, genetic drift and natural selection. Mutations allow the emergence of strains adapted to challenging environments, including various antibiotic types [hauert_evolutionary_2008]. These strains are selected when antibiotics are present in the environment. Critically, the evolution of antibiotic resistance can occur quickly, within days in specific conditions [baym_spatiotemporal_2016], while the development of a new antibiotic typically takes around ten years [wong_estimation_2019]. In this context, it is crucial to understand what conditions favor or hinder the development and spread of antibiotic resistance.
Bacterial populations often have complex spatial structures. For pathogenic bacteria, each patient constitutes a different environment, connected through transmission events. Within a patient, different organs and tissues represent spatial environments between which bacteria can migrate [she_defining_2024]. At a smaller scale, bacteria in the gut microbiota, which can be a reservoir of drug resistance, live in different spatial environments, namely the digesta, the inter-fold regions and the mucus, and some of them form biofilms [donaldson_gut_2016, mccallum_gut_2023, chikina_attheright_2021], which are dense structures of cells enclosed in a polymer-based matrix [costerton_microbial_1995, watnick__biofilm_2000]. Thus, it is important to understand how the spatial structure of microbial populations impacts the evolution and spread of resistance. This question has been explored in epidemiological models in the case of viruses [Debarre07, Althouse13] and of bacteria [krieger_population_2020], and was recently addressed in controlled experiments, for specific antibiotic resistant bacteria [verdon_spatial_2024], but is still relatively under-explored. Spatial structure can lead to small effective population sizes, and small population sizes were recently experimentally shown to substantially impact the evolution and spread of resistance [Coates18, alexander_stochastic_2020]. More generally, the impact of spatial structure on the fate of mutants was studied theoretically [Wright31, Kimura64, maruyama70, maruyama74, Nagylaki80, Barton93, Whitlock97, Nordborg02, whitlock2003, Sjodin05, lieberman_evolutionary_2005, Houchmandzadeh11, allen_evolutionary_2017, marrec_toward_2021, yagoobi_fixation_2021, tkadlec_evolutionary_2023, abbara_frequent_2023] and experimentally [Kryazhimskiy12, Nahum15, france_relationship_2019, chakraborty_experimental_2023, kreger_role_2023], but generally in constant environments.
Here, we ask how spatial structure impacts the fate of a bacterial population suddenly subjected to antibiotic treatment. Is the population eradicated by the drug, or does it survive by developing resistance? We address this question in a minimal model of population structure, composed of identical demes (i.e. subpopulations) connected to each other by identical migrations. This simple structure is known as the island model [Wright31, Kimura64], or the clique or fully-connected graph [lieberman_evolutionary_2005, marrec_toward_2021, abbara_frequent_2023]. We assume that the population initially only comprises drug-sensitive bacteria. Indeed, the initial diversity of a population of pathogenic bacteria in a host is often low, as many infections begin with a small infectious dose [gutierrez_virus_2012, lim_independent_2014, lythgoe_sarscov2_2021, woodward_gastric_2022, hoces_fitness_2023]. For simplicity, we assume that the environment is homogeneous, but note that environmental heterogeneities have a substantial effect on antibiotic resistance  [zhang_acceleration_2011, hermsen_onthe_2012, greulich_mutational_2012], and on evolutionary rescue in structured populations [Uecker14, Tomasini20, Tomasini22]. Specifically, we consider a perfect biostatic drug, which stops the growth of sensitive bacteria, but we also discuss the generalization of our results to other drug modes of action. When a well-mixed population is subjected to a treatment by a perfect biostatic drug (for a long enough time), the population gets extinct, except if resistant mutants are present when the drug is added [marrec_quantifying_2018, marrec_resist_2020, marrec_adapt_2020]. The population can thus be rescued by resistance. We study the impact of population spatial structure on this process. We also analyze how the time when drug is added impacts the fate of a spatially structured population.
We show that spatial structure increases the probability that a bacterial population survives treatment by developing resistance through a neutral mutation. Specifically, for a given time of addition of the antibiotic, the survival probability of the population increases when the migration rate between demes is decreased, and when the number of demes is increased while keeping the same total population size. We show that this is due to the local fixation of resistant mutants in one or several demes. We study the composition of the population versus time before the addition of drug, and find that it is strongly impacted by spatial structure. We further study the time needed for resistant bacteria to colonize all demes once they have fixed in at least one of the demes. Finally, we show that our main results extend to the case where resistant mutants carry a cost, although the timescales are longer. They also extend to the case where resistant mutants are present in the inoculum, and to more complex spatial structures.
Models and methods
Spatially structured populations.
We aim to assess the impact of spatial structure on the establishment of antibiotic resistance, in a minimal model where spatial structure is as simple as possible. Thus, we consider a spatially structured bacterial population comprising $D$ demes (i.e. subpopulations) on the nodes of a clique (i.e. a fully connected graph). This corresponds to the island population model [Wright31]. Each deme has the same carrying capacity $K$ (Figure 1, center). For comparison, we also consider a well-mixed population with the same total carrying capacity $KD$ (Figure 1, left), and a fully subdivided population composed of $D$ demes with carrying capacity $K$ without migrations between them (Figure 1, right). We model migrations between demes through a per capita migration rate $\gamma$, which is the same between all pairs of demes.
Figure 1: Minimal model of spatially structured populations. We consider a clique structure with one deme per node and a per capita migration rate $\gamma$ (center). For large $\gamma$, it becomes a well-mixed population (left). Meanwhile, if $\gamma\to 0$, it becomes a fully subdivided population. Blue and orange markers represent sensitive (S) and resistant (R) bacteria, respectively.
We will also briefly discuss extensions to more complex population structures on graphs [marrec_toward_2021, abbara_frequent_2023].
Bacterial growth and resistance.
We consider two types of bacteria: sensitive (S) and resistant (R) ones. In the absence of antibiotics, the two types of bacteria have fitnesses $f_{S}$ and $f_{R}$, respectively, which set their maximal growth rates (in the exponential phase). They have the same death rate $g$. Thus, we implement selection on birth. However, it is straightforward to generalize our model to selection on death. We assume that growth is logistic in each deme. The growth rate of S (resp. R) individuals is $f_{S}\left(1-N/K\right)$ (resp. $f_{R}\left(1-N/K\right)$), where $N$ is the current population size of the deme considered. When S bacteria divide, their offspring can mutate to R with probability $\mu$. Here, we focus on the rare mutation regime, where $N\mu\ll 1$.
We mainly focus on neutral mutations, but we discuss the impact of a cost or of a benefit of resistance below and in SI Section 3.1. Considering neutral mutations is relevant because resistance mutations are often effectively neutral, e.g. at low drug concentrations. We thus set $f_{S}=f_{R}=1$, implying that the time unit in our system is set by the maximum growth rate of bacteria. We do not model further mutations or backmutations – note that the rate of backmutations was estimated to be one order of magnitude smaller than the rate of compensatory mutations [levin_compensatory_2000].
We consider an ideal biostatic drug that prevents division of the S bacteria, thus changing their fitness to $f_{S}=0$. We assume that R bacteria are not affected by the drug.
Initial conditions and growth regime.
As we aim to model the appearance of resistance, we start with a population comprising only S bacteria. In practice, we initialize each deme (resp. the well-mixed population) with a number of S bacteria representing 10% of $K$ (resp. of $KD$). Then, each deme (resp. the well-mixed population) quickly grows and reaches a steady-state size $K(1-g/f_{S})$ (resp. $KD(1-g/f_{S})$) around which it fluctuates. We choose this initial condition because it is appropriate to model the start of an infection [abel_analysis_2015]. Note that since $N\mu\ll 1$, it is very unlikely that mutations happen during this quick initial growth phase, whose timescale is set by $f_{S}=1$ [marrec_resist_2020]. Our results are thus robust to the initialization, as long as it is performed only with S individuals. We work in the regime where $g\ll f_{S}=1$, so that extinctions of demes are very unlikely in the absence of drug, and demes fluctuate around their steady-state size.
Analytical and numerical methods.
We obtain analytical results from the Moran model [moran_random_1958] and from probability theory. Note that the Moran model assumes that population size is strictly fixed, which is not the case here. However, because we are in the regime where fluctuations around the steady-state size are small in the absence of drug, the Moran process then provides good approximations [marrec_resist_2020].
We perform stochastic simulations of the population evolution, using the Gillespie algorithm [gillespie_general_1976, gillespie_exact_1977]. The details of the simulation framework are described in the SI Section 2.1.
Results
Spatial structure increases the probability that a bacterial population survives treatment
How likely is a spatially structured bacterial population to survive biostatic antibiotic treatment? Because our antibiotic prevents S bacteria from dividing while they have a nonzero death rate, a population is doomed to go extinct in the absence of R bacteria. However, if at least one R individual is present when drug is added, rescue by resistance can happen [marrec_quantifying_2018, marrec_resist_2020]. Here, we aim to assess the impact of spatial structure on rescue by resistance. To this end, we focus on the probability that the bacterial population survives antimicrobial addition in at least one of the demes. Indeed, R bacteria can then spread to the rest of the population via migrations, ensuring its overall survival.
Figure 2A shows the survival probability versus the addition time of the antimicrobial for a spatially structured population with different values of the migration rate. We observe that survival probability increases with drug addition time. Indeed, R mutants are more likely to appear and fix in a population initially composed only by S bacteria if there is more time before drug addition. Moreover, we find that survival probability is higher when migration rate $\gamma$ is smaller.
Figure 2: Survival probability of a structured bacterial population subjected to biostatic antimicrobial. The survival probability is shown versus the drug addition time. Panel A: Spatially structured populations composed of $D=10$ demes with $K=100$ are considered, with different per capita migration rates $\gamma$. Results for the well-mixed population with same total size (“WMP”) are shown for reference. Panel B: Three levels of subdivision of a population with total carrying capacity $KD=1000$, into 5, 10, and 20 demes, are shown for two values of $\gamma$. Results for the well-mixed population (“WMP”) are shown for reference. In all cases, we determined the survival probability from the fraction of $10^{4}$ simulations where the population survived after drug application. The vertical dash-dotted line denotes the average time of appearance of a locally successful mutant in the fastest deme (Equation (2)). The vertical dotted line represents the average appearance time of a successful mutant in a well-mixed population (Equation (1)). Thin lines connecting data points are guides for the eye. The horizontal gray solid line in both panels is the probability of presence of R mutants belonging to a lineage destined to go extinct, calculated from Equation (S16). Parameter values: $f_{S}=1$ without drug, $f_{S}=0$ with drug, $f_{R}=1$, $g=0.1$, $\mu=10^{-5}$.
Figure 2B shows the impact of varying the degree of subdivision of a population of fixed total size, but composed of a different number of demes. We observe that higher subdivision yields higher probabilities of survival.
Local fixation of R mutants promotes the survival of structured populations
To rationalize our results on the impact of spatial structure on the ability of a microbial population to survive the addition of biostatic drug, we focus on rescue by resistance. R mutants can arise in our population with rate $N_{tot}\mu g$, where $N_{tot}$ is the total number of individuals in the population, $\mu$ the mutation probability upon division and $g$ the death rate, which equals the division rate at steady state. All the populations we considered in Figure 2 have the same total steady-state size $N_{tot}=KD(1-g/f_{S})$ in the absence of drug. Thus, the mutant supply is not impacted by spatial structure. However, the fate of mutants can be impacted by spatial structure. Most R mutants that appear give rise to lineages that go extinct, whatever the structure. The survival time of their lineages is not impacted by population size. Let us now consider mutants that fix. Since we consider neutral mutants, their fixation probability is $1/N$ in a population of size $N$. In spatially structured populations, fixation can occur locally within a deme, and then it is governed by the smaller steady-state population size $N=K(1-g/f_{S})$ of a deme. If migrations are rare enough, once R individuals have fixed in a deme, they survive there, and can rescue the population.
To understand the impact of local mutant fixation on population survival, let us focus on successful mutants, i.e. those that give rise to a lineage that fixes. The average time of appearance $\langle t_{af\,W}\rangle$ of a successful mutant in a well-mixed population of size $N$ is given by:
$\langle t_{af\,W}\rangle=t_{app}\times\frac{1}{p_{fix}}=\frac{1}{N\mu g}\times N=\frac{1}{\mu g}\,,$ (1)
where $t_{app}=1/(N\mu g)$ is the average appearance time of a mutant and $p_{fix}=1/N$ is its fixation probability. Recall that we are considering neutral mutants (see above, and SI Section 1.1). Importantly, this result does not depend on population size $N$. Therefore, it holds both for the well-mixed population and for each deme in a structured population.
In a structured population, the first mutant that fixes locally in one deme may rescue the population by its resistance. The appearance of such a locally successful mutant is a Poisson process with rate $\mu g$ (see Equation (1)). Thus, the average appearance time of a locally successful mutant in the fastest deme is (see SI Section 1 for details):
$\langle t_{af\,F}\rangle=\frac{1}{D\mu g}\,.$ (2)
Crucially, this is $D$ times faster than the appearance of a successful mutant in a well-mixed population (which is given by Equation (1)).
The addition of a biostatic antimicrobial (leading to $f_{S}=0$, $f_{R}=1$ when drug is added) should have a different effect depending how the addition time $t_{add}$ compares to the two timescales defined in Equations (1) and (2). Specifically, we predict that:
  • If $t_{add}\ll\langle t_{af\,F}\rangle$, the bacterial population is likely to be eradicated, whatever its structure.
  • If $\langle t_{af\,F}\rangle\ll t_{add}\ll\langle t_{af\,W}\rangle$, it is likely that a locally successful mutant has appeared. This rescues a structured population. However, it is likely that no successful mutant has appeared in the well-mixed population yet. A well-mixed population is thus expected to go extinct.
  • If $t_{add}\gg\langle t_{af\,W}\rangle$, both the well-mixed and the structured population are expected to survive.
In Figure 2, we show the two timescales defined in Equations (1) and (2). We observe that the most substantial impact of structure on survival probabilities is observed when $\langle t_{af\,F}\rangle<t_{add}<\langle t_{af\,W}\rangle$. This is fully in line with our theoretical analysis based on timescale comparisons. Furthermore, we observe that all populations survive if $t_{add}\gg\langle t_{af\,W}\rangle$. Finally, if $t_{add}\ll\langle t_{af\,F}\rangle$, the population is eradicated in most cases. It can nevertheless be rescued by non-successful mutant lineages (see also Figure S3). Thus, the probability of treatment survival is given by the probability that non-successful mutants are present when the drug is added, see Equation (S16). Our simulation results are in good agreement with this analytical prediction, see Figure 2.
Spatial structure impacts population composition before drug addition
To understand in more detail the role of spatial structure on resistance spread, let us study the system composition before drug is added.
Figure 3A shows the time evolution of the average number of R mutants present in the whole structured population. We observe that it is not impacted by population structure. Whatever the migration rate, the number of R individuals grows as mutants fix in some replicates of our stochastic simulations. This occurs either first locally for structured populations, or directly in the whole population for the well-mixed population (see SI Section 3.3), but the average over replicates remains the same. The long-time limit corresponds to the total steady-state population size, $N=K\left(1-g/f_{R}\right)$, as R mutants take over the whole population. This occurs with a timescale of order $\langle t_{af\,W}\rangle$. Figure 3B shows that the variance across replicates of the number of R mutants is strongly impacted by population structure. This is particularly true in the time interval between $\langle t_{af\,F}\rangle$ and $\langle t_{af\,W}\rangle$. This suggests that the impact of population structure on the variance of mutant number is driven by R mutant fixation. Figure 3B further shows that the variance of the number of mutants across replicates is larger when migrations are more frequent. This can be understood qualitatively by comparing a well-mixed population and a structured population with small migration rate. In the first case, mutants fix in one step, while in the second one they fix separately in each deme (see Figure S2). The large variance observed for the well-mixed population comes from the variability across replicates of the appearance time of a successful mutant. For the structured population with small migration rate, this is partly smoothed by lumping together $D$ separate demes with different fixation times. Moreover, the appearance time of a successful mutant in the fastest deme has a smaller variance than that in a well-mixed population (see SI Section 3.3). Here, we focused on mutant fraction in the whole population because we are interested in the overall behavior of the population, but the dynamics at the deme level is described in SI Section 3 and below in Figure 4.
Figure 3: Impact of spatial structure on population composition in the absence of drug. Panel A: The mean across replicates of the total number of R mutants in the population is shown versus time for population structures differing by the migration rate $\gamma$, and for the well-mixed population with same total size (“WMP”). Horizontal dashed line: steady-state total population size. Panel B: Variance across replicates in the total number of R mutants in the population, in the same conditions as in panel A. Vertical dotted line in panels A and B: time when largest variance was obtained in panel B (t$=5\times 10^{5}$). Panel C: Population composition at t$=5\times 10^{5}$, for different migration rates. Results are obtained from $10^{4}$ replicate simulations. Parameter values (in all panels): $K=100$, $D=10$, $f_{S}=f_{R}=1$, $g=0.1$, $\mu=10^{-5}$.
To go beyond the mean and variance in overall mutant number, in Figure 3C we examine the population composition in the absence of drug, at the time when we obtained the largest variance in Figure 3B (vertical dotted line). We describe the population composition by defining four categories: either there are no R bacteria at all (“No R” - red), or at least one deme comprises between 1 and 9 R bacteria but no deme has more than 9 R (“Small R population” - yellow), or at least one deme comprises more than 9 R individuals, but the R type has not fixed in the whole population (“Big R population” - cyan), or R bacteria have fixed in the whole population (“Only R” - blue). Note that we distinguish “Small R population” from “Big R population” to account for the possibility of stochastic extinction of R lineages, which becomes negligible if at least 10 R bacteria are present in a deme (see SI Section 3). In Figure 3C, we observe that in most replicates, at the time considered, the well-mixed population has either fixed resistance or does not have any mutants. Thus, the number of mutants is either 0 or $(1-g/f_{R})KD$, yielding the large variance mentioned above. As we partition the system spatially and lower the migration rate, the composition of structured populations becomes more mixed, and the fraction of replicates with overall fixation or overall extinction decreases. This results in a smaller variance across replicates, as the number of mutants becomes more homogeneous across them.
An important cause of the composition difference between structured and well-mixed populations is the possibility of local fixation of R mutants. What is the fixation dynamics at the deme level before we add the drug? In Figure 4, we report how many demes have fixed resistance for different migration rates and at different times, in the form of histograms computed over simulation replicates. Note that for the well-mixed population we report overall fixation. As expected, the number of demes where R mutants have fixed increases over time. Furthermore, we observe that the distribution of demes that have fixed resistance and its dynamics are strongly impacted by spatial structure. For large migration rates (panel A), at all times, most replicates have fixed resistance in either no deme or all demes. This is close to the large-$\gamma$ limit (i.e. the well-mixed population), shown as horizontal lines in panel A. When the migration rate $\gamma$ is decreased (panels B and C), the fraction of realizations featuring an intermediate number of demes that have fixed resistance increases at intermediate times. For the small migration rate $\gamma=10^{-7}$ (panel C), the distributions become close to those observed for $\gamma=0$ (panel D). Thus, small migration rates result in a population with a transient strong heterogeneity across demes. This has a crucial impact on the outcome of drug treatment, and on the survival probability curves in Figure 2. Indeed, the biostatic drug does not affect R mutants, and the presence of at least one deme where R has fixed rescues the population when the drug is applied.
Figure 4: Number of demes that have fixed resistance before drug application. Histograms of the number of demes that have fixed resistance, computed across $10^{4}$ simulation replicates, are shown for different values of the elapsed time t (on different curves with different colors), and of the migration rate $\gamma$ (in different panels). Specifically, for each $i\in[0,10]$, we show the fraction of realizations in which $i$ demes have fixed resistance. The horizontal line markers in panel A show the results for a well-mixed population with same total size (“WMP”). Parameter values: $K=100$, $D=10$, $f_{S}=f_{R}=1$, $g=0.1$, $\mu=10^{-5}$.
R mutants readily colonize the whole population after drug addition
So far, we focused on the impact of spatial structure on the survival probability of the population, and on its composition before the addition of drug. A key conclusion is that local fixation of R mutants allows earlier population survival. Let us now ask what happens after drug addition. Since our biostatic drug prevents S bacteria from dividing while mutations occur upon division, there are no new mutations after drug addition. If R mutants fixed in at least one deme before drug addition, how fast do they spread in all demes of a spatially structured population?
To address this question, let us first calculate the time $\tau_{c\,mig}$ it takes for an R mutant to migrate to a deme without R mutants, starting from a population where resistance has fixed in $k\geq 1$ demes out of $D$. Then, there are $Nk$ mutant individuals, where $N$ is the steady-state deme size, which can migrate to each of the $D-k$ uncolonized demes at a per capita rate $\gamma$. However, once an R mutant arrives in a new deme, its lineage may go stochastically extinct. This happens with probability $1-g/f_{R}$, assuming that the deme does not contain other individuals (see derivation in SI Section 1.5). Thus, we have:
$\tau_{c\,mig}(k)=\frac{1}{\gamma Nk(D-k)}\frac{1}{1-g/f_{R}}=\frac{1}{\gamma Nk(D-k)}\,C\,,$ (3)
where $C=1/(1-g/f_{R})$ captures the effect of stochastic extinctions. In Figure 5A, we compare the analytical prediction in Equation (3) to our simulation results, obtaining a good agreement. We also show the impact of stochastic extinctions by setting $C=1$ in the analytical predictions: their effect is not negligible here (of order 10%, and they would be larger if $g$ was closer to $f_{R}$). The time $\tau_{c\,mig}$ for R mutants to colonize the next deme features a minimum when half of the demes are mutant. A larger proportion of migration events then leads to colonization. Indeed, more mutant demes yield more mutants that may migrate, but fewer wild-type demes where they can fix, leading to a trade-off.
Equation (3) neglects the presence of S bacteria in the population, which would affect the stochastic extinction of R lineages. This is acceptable provided that the extinction of S bacteria upon drug addition is fast enough. The decay time $\tau_{S}$ of a well-mixed population comprising $N$ S bacteria upon drug addition reads [marrec_resist_2020]:
$\tau_{S}=\frac{1}{g}\sum_{i=1}^{N}\frac{1}{i}\,.$ (4)
For the migration rate considered in Figure 5A, we checked that $\tau_{S}$ is indeed negligible with respect to $\tau_{c\,mig}$.
Figure 5: Colonization of the whole population by R mutants. Panel A: The time $\tau_{c\,mig}$ to colonization of the next deme by R mutants through migrations is shown versus the number $k$ of demes where R mutants have already fixed. Migration rate: $\gamma=10^{-6}$. We compare simulation data from $5\times 10^{3}$ realizations with the theoretical prediction from Equation (3), accounting for stochastic extinctions (“Theory”) or not (“Theory, $C=1$”). Error bars represent 95% CI. Panel B: The average time to complete colonization (i.e., when population size first becomes at least equal to $0.9K$ in each separate deme) is shown versus the drug addition time $t_{add}$ for structured populations with different migration rates, and for the well-mixed population with same total size (“WMP”). Out of $10^{4}$ realizations, we restrict to simulations where the population survived treatment. Solid black line: $y=x$ diagonal. Horizontal lines: $t_{c\,tot}(\gamma)$ for the two lowest migration rates $\gamma$. Vertical lines: average local fixation time of R mutants in the fastest deme ($\langle t_{af\,F}\rangle$), in a well-mixed population ($\langle t_{af\,W}\rangle$), and in the slowest deme ($\langle t_{af\,S}\rangle$). Confidence intervals are too small to be visible in this logarithmic scale. Parameter values for both panels: $K=100$, $D=10$, $f_{S}=f_{R}=1$, $g=0.1$, $\mu=10^{-5}$.
So far, we focused on one step of the spread of R mutants. How long does it take in practice for R mutants to colonize the whole structured population? In Figure 5B, we show the total colonization time versus drug addition time for different migration rates. We observe different regimes in these curves depending on the drug addition time $t_{add}$.
First, if the drug is added after a short time, at most one deme in the population will comprise R mutants or will have fixed them. This should happen if the drug addition time $t_{add}$ is shorter than the average appearance time $\langle t_{af\,F}\rangle$ of a locally successful mutant in the fastest deme. How long does it then take for R mutants to colonize the whole population? To obtain this total colonization time $t_{c\,tot}$, we can sum $\tau_{c\,mig}$ over the $D-1$ steps needed to sequentially colonize the structured system. It is thus given by:
$t_{c\,tot}=\sum_{k=1}^{D-1}\tau_{c\,mig}(k)=\frac{2(\Gamma+\psi(D))}{D\gamma N}\frac{1}{1-g/f_{R}}\,,$ (5)
where $\Gamma$ is the Euler gamma constant, while $\psi$ is the digamma function. In Figure 5B, we report the overall colonization time of the population by R mutants, defined as the time $t_{\textrm{add}}$ when drug is added plus the time needed for colonization after drug addition. We observe that for small addition times satisfying both $t_{add}\ll\langle t_{af\,F}\rangle$ and $t_{add}\ll t_{c\,tot}$, $t_{c\,tot}$ indeed describes well the colonization time. Indeed, it is then unlikely that mutants are present in more than one deme when drug is added, and the addition time has a negligible contribution. It leads to a plateau whose value is governed by the migration rate. Note that in Figure 5B, this plateau is only observed for small migration rates ($\gamma\lesssim 10^{-5}$) such that the condition $t_{add}\ll t_{c\,tot}$ is satisfied for the smallest $t_{add}$ values considered. For larger values of $t_{add}$ and larger migration rates, the timescale $t_{add}$ dominates the overall colonization time of the population by R mutants, as we have $t_{add}>t_{c\,tot}$. In these cases, mutant spread can be considered fast when drug is added (see Figure 5B). Finally, if $t_{add}$ is further increased, overall colonization of the population by R mutants may happen before drug addition. Thus, the overall colonization time becomes smaller than $t_{add}$ and tends toward a new plateau. Note that in this regime, colonization can occur before drug addition through migrations or new mutations, followed by fixations. The time it takes for a locally successful mutant to appear in the slowest deme can be expressed as $\langle t_{af\,S}\rangle=\langle t_{af\,W}\rangle\sum_{i=1}^{D}1/i$ (see SI Section 1.2). It is beyond this time that we expect a plateau, even for small migration rates. This is indeed what is observed in Figure 5B.
Extension to cases where resistance carries a cost or a benefit
So far, we considered that the resistant mutant is neutral. However, antibiotic resistance often carries a cost in the absence of drug [Andersson99, Andersson10]. Besides, if antibiotic is already present at relatively low doses in the environment before the addition of drug, it can induce a benefit of resistance before drug is added. Let us briefly consider these cases.
First, our results extend to effectively neutral mutants having a fitness cost $\delta$ such that $N\delta\ll 1$. The impact of structure should in fact be even stronger if $N\delta\ll 1$ but $ND\delta\gg 1$ because fixation is then strongly suppressed in a well-mixed population of size $ND$. Second, the case where the cost of resistance $\delta$ is stronger ($N\delta\gg 1$) is discussed in SI Section 3.1. While spatial structure should have a strong effect here, this will only matter for the very long times associated to the fixation of deleterious mutants. Third, the effect of spatial structure we evidenced here for neutral mutants does not extend to the case where resistant mutants are beneficial, see SI Section 3.1.
Extension to an inoculum that contains R mutants
So far, we focused on a sensitive inoculum and on resistant individuals appearing through mutations. The case where mutants are already present in the inoculum is also interesting. We consider it in SI Section 5. Our main results about the increase in probability of treatment survival granted by a spatial population structure extends to the case where R mutants are present in the inoculum (see also Ref. [kreger_role_2023]). Specifically, we observe that in this case too, the survival probability of the population increases when the mutation rate decreases, see Figure S6.
More complex spatial structures behave similarly to the clique
So far, we considered a minimal model of spatially structured population, known as the clique, where all demes are equivalent and connected to one another by identical migration rates. Do our findings extend to more complex spatial structures? To address this question, we explored two structures with more reduced symmetry.
First, we consider a two-dimensional square lattice with periodic boundary conditions. Here, each deme is directly connected to its four nearest neighbors, thus only allowing for local migrations. We compare the lattice to the clique with the same total outgoing migration rate from a deme. Our simulations yield no statistically significant difference for the survival probability in the square lattice and in the clique (see SI Figure S5). This suggests that making migrations more local does not affect treatment survival.
Second, we consider a star structure, comprising a central deme connected to $D-1$ leaves [marrec_toward_2021]. All leaves are assumed to be equivalent. Denoting by $\gamma_{I}$ (resp. $\gamma_{O}$) the migration from a leaf to the center (resp. from the center to a leaf), we defined migration asymmetry as $\alpha=\gamma_{I}/\gamma_{O}$. We compared the star with a given outgoing migration rate from a leaf to the clique with the same total outgoing migration rate from a deme (i.e. $\gamma_{0}=(D-1)\gamma$). We observe in Figure S5 that the star gives the same survival probability to treatment as the clique and the lattice if $t_{add}<\langle t_{af\,F}\rangle$, i.e., in cases where usually not more than one deme has fixed resistance before drug addition. For larger values of drug addition time $t_{add}$, our numerical results for the star depend on $\alpha$ and significantly differ from those obtained for the clique and the lattice, although the difference remains minor. We interpret this minor difference as arising from wild-type individuals possibly re-invading demes where mutants have fixed, which is structure-dependent. See SI Section 4 for more details. Overall, our main results are robust to changing the graph on which demes are placed.
Discussion
Here, we showed that spatial structure can facilitate the survival of a bacterial population to antibiotic treatment, starting from a sensitive inoculum. Indeed, the bacterial population can be rescued if antibiotic resistant mutants are present when drug is added. While the emergence of resistant bacteria by random mutations only depends on total population size and not on spatial structure, their fate can be affected by spatial structure. In the relevant case where the mutation that provides resistance is neutral or effectively neutral, its probability of fixation is increased in smaller populations. This leads to local fixation of resistant mutants in demes, which then constitute refugia of resistance mutants and allow the population to survive when drug is added. Because of this, spatial structure facilitates evolutionary rescue by cost-free drug resistance. The survival probability of the population increases when the migration rate between demes is decreased, and when the number of demes is increased. Once the population is rescued by resistance, migrations allow resistant mutants to spread in all demes. While our main results were obtained starting from a sensitive inoculum, we showed that they extend to the case where there are resistant mutants in the inoculum. While we considered a minimal model where all demes are equivalent and connected to each other with identical migration rates, bacterial populations can have diverse spatial structures, with migrations that may be more local or asymmetric. We showed that our main conclusions still hold for a lattice of demes, modeling more localized migrations, and for demes located on a star graph with asymmetric migrations.
The effect of spatial structure we evidenced here is due to stochastic fixation of neutral mutants. Stochastic effects often give rise to original behaviors of small and structured populations. In particular, inoculum size impacts the survival of bacterial population in the presence of antibiotic [Coates18, alexander_stochastic_2020], and spatial partitioning impacts the efficiency of antibiotic resistance by production of the beta-lactamase enzyme [verdon_spatial_2024]. However, the effect evidenced here is different, as it relies on the stochasticity in the fate of a mutant, and not in the stochasticity in the inoculum size.
Because local fixation of neutral resistant mutants is instrumental to our effect, it involves rather long timescales. The average appearance time $\langle t_{af\,F}\rangle=1/(D\mu g)$ of a locally successful mutant in the fastest deme is of particular relevance. As it scales as the inverse mutation probability per individual and per generation $1/\mu$, it can be quite long, even considering that there are often several different mutational targets that give rise to resistance, leading to a larger effective $\mu$. However, this timescale is inversely proportional to the number $D$ of demes, meaning that it can potentially become arbitrarily small when increasing population subdivision. With a mutation probability per nucleotide and per generation of $10^{-10}$ [lee_rate_2012], assuming 10 possible mutations for the development of resistance to a given drug, gives $\mu=10^{-9}$. Then, for bacteria dividing once per hour, we find $\langle t_{af\,F}\rangle=10^{8}$ h for $D=10$, but $\langle t_{af\,F}\rangle=10^{2}$ h for $D=10^{7}$. Such large numbers of demes are realistic for instance in the case of intestinal glands [Nguyen10]. Note also that larger mutation rates, e.g. in mutators, can be relevant to the evolution of resistance [Gifford23]. While we mainly focused on neutral mutants giving resistance, our findings extend to effectively neutral mutants, and to deleterious mutants. However, in the latter case, the timescales involved would be even longer.
Here, we modeled the action of the antibiotic as preventing any division of sensitive bacteria. In other words, we considered a perfect biostatic drug. However, our main findings generalize to biocidal drugs that increase the death rate of sensitive bacteria, or to drugs which combine both modes of actions. Indeed, what happens before the addition of drug, in particular the local fixation of resistant mutants, is not impacted by the mode of action of the drug. Thus, the effect of spatial structure we evidenced here holds independently of this. The difference between the biostatic and the biocidal case lies in the decay of the sensitive bacteria once drug is added. In the biocidal case, sensitive microbes may still divide during this phase, and new resistant mutants may then appear [marrec_resist_2020, czuppon_stochastic_2023]. This is the case both in well-mixed and in structured populations, with only minor differences between them, due to the stochasticity of extinction. Thus, our main findings are robust to the mode of action of the drug. Beyond modes of action, it would be interesting to study the impact of spatial structure on multi-step drug resistance evolution [Igler21].
Here, we considered a minimal model of spatial structure, without any environmental heterogeneities. This allowed us to find a generic effect of spatial structure on the survival of a population to antibiotic treatment. It would be very interesting to extend our work to heterogeneous environments, which are known to have a substantial effect on antibiotic resistance evolution [zhang_acceleration_2011, hermsen_onthe_2012, greulich_mutational_2012], and on evolutionary rescue in structured populations [Uecker14, Tomasini20, Tomasini22].
Some mechanisms of resistance to antibiotics involve the production of a public good. One prominent example is the beta-lactamase enzyme, which degrades beta-lactam antibiotics in the environment. Spatial structure was recently experimentally shown to have an important impact in this case [verdon_spatial_2024]. Note that collective protection via a public good exists in multiple other cases [yurtsev_oscillatory_2016, domingues_social_2017, galdino_sideropores_2024]. Recently, a model was developed to describe this phenomenon in a well-mixed population [hernandez_coupled_2023, hernandez_ecoevolutionary_2024]. Extending this study to the case of structured populations would be very interesting. More generally, coupling these evolutionary questions to ecological interactions, which also have an interesting interplay with spatial structure [limdi_asymmetric_2018, gokhale_migration_2018, wu_modulation_2022], is an exciting perspective.
It would be interesting to test our predictions experimentally. In vitro experiments considering bacterial populations with migrations between demes have been performed [Kryazhimskiy12, Nahum15, france_relationship_2019, chakraborty_experimental_2023, kreger_role_2023], but generally with a constant environment. Investigating the fate of populations upon the addition of drug in these setups would be very interesting. A challenge is that the timescales relevant here are long. However, they can become smaller if the number $D$ of demes is increased. Long experiments with large numbers of demes can be achieved using robots for serial passage [Kryazhimskiy12, Horinouchi14, Maeda20], and the connection between theory and experiment for spatially structured populations is progressing [abbaraPreprint].
Code availability
Python code for our stochastic simulations is freely available at https://github.com/Bitbol-Lab/DrugRes_StructPop
Acknowledgments
This research was partly funded by the Swiss National Science Foundation (SNSF) (grant No. 315230_ 208196, to A.-F. B.) and by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 851173, to A.-F. B.).
Supplementary information
1 Derivation of analytical results presented in the main text
1.1 Appearance of a mutant that fixes in a well-mixed population
In our model, mutations from sensitive to resistant individuals happen at birth at per capita rate $\mu$. Thus, since at equilibrium $f_{S}\left(1-N/K\right)=g$, the average time it takes for a mutant to appear in a population of $N$ wild-types is:
$t_{app}=\frac{1}{N\mu g}.$ (S1)
The fixation probability of neutral mutants in a population of size $N$ is:
$p_{fix}=\frac{1}{N}.$ (S2)
Thus, the average time of appearance of a mutant that fixes in a well-mixed population is given by:
$\langle t_{af\,W}\rangle=t_{app}\times\frac{1}{p_{fix}}=\frac{1}{\mu g}.$ (S3)
1.2 Appearance of the slowest mutant that fixes in a structured population
Let us now focus on the appearance of a mutant that fixes in the slowest deme in a population composed of $D$ independent demes, in the absence of migrations. By slowest deme, we mean the deme in which a locally successful mutant (i.e. a mutant destined to fix locally in its deme of origin) takes the longest time to appear.
For a mutant to fix in the slowest deme, mutants must have fixed in all others. Let us denote by $t_{af\,S}$ (resp. $t_{af\,d}$) the time of appearance of a mutant in the slowest deme (resp. in any deme). The probability $P(t_{af\,S}\leq t)$ that $t_{af\,S}$ is smaller or equal than $t$ reads for any $t$:
$P(t_{af\,S}\leq t)=[P(t_{af\,d}\leq t)]^{D}.$ (S4)
The appearance of a locally successful mutant is a Poisson process with rate $\mu g$ (corresponding to the inverse of Equation (S3)). Thus, the time $t_{af\,d}$ is exponentially distributed with rate $\mu g$ and probability density
$p_{d}(t_{af\,d}=t)=\mu ge^{-\mu gt}\,,$ (S5)
and we have $P(t_{af\,d}\leq t)=1-e^{-\mu gt}$. This allows us to express the probability density $p_{S}(t)$ of appearance of a mutant in the slowest deme:
$\begin{split}p_{S}(t)dt&=dp(t_{af\,S}\in[t,t+dt])=P(t_{af\,S}\leq t+dt)-P(t_{af\,S}\leq t)\\ &=\frac{dP(t_{af\,S}\leq t)}{dt}dt=D\left[P(t_{af\,d}\leq t)\right]^{D-1}\,\frac{dP(t_{af\,d}\leq t)}{dt}dt\\ &=D\left[1-e^{-\mu gt}\right]^{D-1}\mu ge^{-\mu gt}dt\,.\end{split}$ (S6)
Note that $p_{S}(t)$ is positive for all $t$ and normalized, as expected.
The average appearance time of a locally successful mutant in the slowest deme is thus given by:
[…]
which involves the harmonic number $\sum_{i=1}^{D}1/i$.
1.3 Appearance of the fastest mutant that fixes in a structured population
Let us now consider the fastest deme, i.e. the deme where a locally successful mutant first appears. By definition, if a successful mutant takes longer than $t$ to appear in the fastest deme, it entails that successful mutants will take longer to appear in all demes. Denoting by $t_{af\,F}$ the time of appearance of a successful mutant in the fastest deme, we have:
$P(t_{af\,F}>t)=[P(t_{af\,d}>t)]^{D}.$ (S8)
This allows us to express the probability density $p_{F}$ of appearance of a successful mutant in the fastest deme:
$\begin{split}p_{F}(t)dt&=dp(t_{af\,F}\in[t,t+dt])=-\frac{dP(t_{af\,F}>t)}{dt}dt\\ &=-D[P(t_{af\,d}>t)]^{D-1}\frac{dP(t_{af\,d}>t)}{dt}dt\\ &=-D[1-P(t_{af\,d}\leq t)]^{D-1}\left[-\frac{dP(t_{af\,d}\leq t)}{dt}\right]dt\\ &=D\mu ge^{-\mu gtD}dt\,.\end{split}$ (S9)
Note that $p_{F}(t)$ is positive for all $t$ and normalized, as expected.
The average appearance time of a successful mutant in the fastest deme is thus given by (see also Ref. [bitbol_quantifying_2014]):
[…]
1.4 Survival probability of the population with $\gamma=0$
When there is no migration between demes, i.e. when $\gamma=0$, the survival probability $p_{s\,\gamma=0}(D)$ of a population comprising $D$ demes, is related to the survival probability $p_{s\,d}$ of a deme through:
$p_{s\,\gamma=0}(D)=1-(1-p_{s\,d})^{D}.$ (S11)
Indeed, denoting by $p_{e\,\gamma=0}$ the probability of extinction of all demes, we have:
$p_{s\,\gamma=0}(D)=1-p_{e\,\gamma=0}(D).$ (S12)
Let $p_{e\,d}$ be the probability that the population in one deme becomes extinct. When $\gamma=0$, extinctions in different demes are independent events with the same probability, thus:
$p_{s\,\gamma=0}(D)=1-p_{e\,\gamma=0}(D)=1-\left(p_{e\,d}\right)^{D}=1-(1-p_{s\,d})^{D}\,.$ (S13)
We checked that Equation (S11) was satisfied in our numerical simulations, within errors given by the standard error of the proportion.
1.5 Correction for stochastic extinction in the mutant spreading time
We derive here the correction accounting for the stochastic extinction of mutants, which was used in the main text to obtain the mutant colonization time in Equation (3), namely $C=(1-g/f_{R})^{-1}$. Let us consider one resistant bacterium with replication rate $f_{R}$ and death rate $g$. Let us denote by $e$. the extinction probability of its lineage. This resistant individual can die without dividing with probability $p_{D}=g/(f_{R}+g)$, giving an extinction probability is $e_{D}=1$. It can also divide before dying with probability $p_{R}=f_{R}/(f_{R}+g)$, leading to the extinction probability $e_{R}=e^{2}$. We can then write an equation for the extinction probability $e$ of the lineage of the single resistant bacterium:
$e=p_{D}\,e_{D}+p_{R}\,e_{R}=\frac{g}{f_{R}+g}+e^{2}\frac{f_{R}}{f_{R}+g}.$ (S14)
The solutions to this equation are $e=1$ or $e=g/f_{R}$ if $g<f_{R}$. Then, the survival probability reads: $1-g/f_{R}$ if $g<f_{R}$. The correction for stochastic extinction in the mutant colonization time (Equation (3)) is then a multiplicative factor $C=(1-g/f_{R})^{-1}$.
1.6 Probability of presence of mutants from a lineage destined to go extinct
Let us consider a deme where R mutants have not fixed, in the absence of drug. Let us determine the probability $p_{pres}$ of presence of mutants that arise through mutations upon division of sensitive individuals, and whose lineage is destined to go extinct. We follow the method exposed in Ref. [marrec_resist_2020], which is based on the transition rate matrix of the process.
We distinguish each case where exactly $i$ mutants are present. In each case, the contribution to $p_{pres}$ is obtained by multiplying the probability $p_{R}$ that at least one R mutant is present, by the conditional probability $p_{R}^{c}(i)$ of having $i$ R mutants, provided that a lineage of mutants exists. We have $p_{R}=\tau_{R}^{d}\,N\mu g$, where $\tau_{R}^{d}$ is the the average lifetime of a resistant lineage destined to extinction, and $N\mu g$ the total mutation rate in the deme, with $N$ the steady-state deme size, $g$ the death rate, $\mu$ the mutation rate. Meanwhile, $p_{R}^{c}(i)=\tau_{R,i}^{d}/\tau_{R}^{d}$, with $\tau_{R,i}^{d}$ the average time the lineage spends in a state with $i$ mutants, calculated as $\tau_{R,i}^{d}=-\pi_{i}/\pi_{1}(\mathbf{\tilde{R}}^{-1})_{i1}$. Here, $\pi_{i}$ is the probability that resistant mutants go extinct, starting from $i$ of them, while $\mathbf{\tilde{R}}$ is the reduced transition rate matrix for mutants (i.e., the transition rate matrix where rows and columns corresponding to absorbing states are eliminated [marrec_resist_2020]).
Summing the term $p_{R}\,p_{R}^{c}(i)$ over all possible $i$, we obtain the probability $p_{pres}$ that mutants from a lineage destined to go extinct are present in the deme:
$p_{pres}=-\sum_{i=1}^{N}Ng\mu\,\frac{\pi_{i}}{\pi_{1}}(\tilde{\mathbf{R}}^{-1})_{i1}\,.$ (S15)
As discussed in the main text, when drug is added to the system, the population can survive if mutants are present when the drug is added. For survival, it is also necessary that the mutant lineages do not go stochastically extinct. Thus, combining $p_{pres}$ with the stochastic extinction probability (see Subsection 1.5) yields an estimate of the survival probability $p_{surv}$ assuming that mutants have not fixed:
$p_{surv}=-\sum_{i=1}^{N}\left[1-(g/f_{R})^{i}\right]\,Ng\mu\,\frac{\pi_{i}}{\pi_{1}}(\tilde{\mathbf{R}}^{-1})_{i1}\,.$ (S16)
2 Description of the simulations performed in the main text
2.1 General simulation approach
Our simulations are based on the Gillespie algorithm [gillespie_exact_1977, gillespie_general_1976, higham_modeling_2008]. Let us consider a clique population structure composed of $D$ demes, labeled with $i\in[1,D]$. Let $S_{i}$ (resp. $R_{i}$) denote one sensitive (resp. resistant) individual in deme $i$. The system obeys the following reaction network:
[…]
The action of the biostatic drug is modeled through a change of the fitness of sensitive individuals, $f_{S}$, as follows:
$f_{S}=\begin{cases}1\hskip 8.5359pt\text{in the absence of drug,}\\ 0\hskip 8.5359pt\text{after biostatic drug addition.}\end{cases}$ (S18)
2.2 Simulation runs and analyses
In general, we run simulations as follows. For each drug addition time $t_{add}$ considered, we run multiple simulation replicates. In each of them, we initialize the system with $0.1\times K$ sensitive individuals in each deme (or $0.1\times DK$ in the well-mixed population), let the system evolve until $t_{add}$ and save its state. Then we add the biostatic drug, and we let the system evolve until either extinction of the population or full system colonization by resistant individuals. If resistance fixes before drug addition, we save the state at this point and stop the simulation. When stopping the simulation, in all cases described above, we save the stopping time. Because of stochastic fluctuations around the mean, we consider that mutants have colonized one deme (resp. the whole population in the well-mixed case) if the population size is greater than $0.9\times K$ (resp. $0.9\times DK$).
Below, we describe in more detail how simulations are performed and analyzed for each of the main figures.
Survival probability.
In Figure 2, we calculate the survival probability as the fraction of simulation replicates where bacteria survive the application of the drug, in the sense that one deme comprises at least $0.9\times K$ mutants. Note that this allows us to exclude cases in which mutants quickly go stochastically extinct after drug addition. In the structured populations, survival is intended to be in at least one of the demes.
Average and variance of mutant numbers at the population level.
In Figure 3A-B, for each time and each structure, we perform simulation replicates in the absence of drug. For each of these replicates, we count the number of mutants in the population. We then calculate the average and variance of the total number of mutants across the different replicates.
Average and variance of mutant numbers at the deme level.
In Figure S4, for each time and each structure, we perform simulation replicates in the absence of drug. For each replicate we choose one deme (deme 1 in practice in Figure S4, but the result is similar for other demes given the symmetry of the clique). We calculate the average and variance of the number of mutants in that deme across stochastic realizations.
Population composition without drug.
In Figure 3C, we analyze the composition of the population before the addition of drug. In the state of the system saved before environmental change, we focus on the number of mutants, and partition the outcomes of our stochastic simulations into the four categories described in the main text (“Only R”, “Big R pop.”, “Small R pop.”, “No R”). We report the fraction of replicates falling in each of these categories.
Number of demes where resistance has fixed.
In Figure 4, for each time and a structure, we perform multiple replicates. For each of them, we calculate the number of demes where the mutant type has fixed before drug addition. We report a histogram of these number of demes across replicates.
Time until colonization of next deme.
In Figure 5A, we perform simulations to assess the time until colonization of next deme. These simulation are set up slightly differently than others, to focus on the colonization process. We initialize one of the demes with $N=0.9\times K$ mutants, all others with $N=0.9\times K$ wild-types, mimicking fixation of mutants in one deme before drug addition. We add the biostatic drug at $t=0$ in this simulation, thus modeling the case where one deme has fixed resistance before drug addition. We then let the system evolve until overall colonization of the structured population by resistant mutants. For each $k\geq 1$, we save the times of successful colonization of the $(k+1)^{th}$ deme, given that $k$ demes were already colonized by mutants. We define successful colonization of a deme as reaching a number of mutants $0.9\times K$. We then obtain the values of $\tau_{c\,mig}(k)$ as the differences between the successive times we recorded.
Time until overall colonization.
In Figure 5B, we report the time until overall colonization of the structured population by resistant mutants. When saving the state of the system after overall mutant colonization (resp. extinction), we also record the time it took for overall colonization (resp. extinction). Thus, we can calculate the mean of the overall colonization time for each $t_{add}$ and each structure. Note that the fixation of resistance may happen before drug addition (this happens in particular for addition times longer than $\langle t_{af\,S}\rangle$). In this case, we consider that colonization happens when resistant individuals reach fixation in all demes and there are at least $0.9\times K$ resistant individuals.
3 Clique population structure with sensitive inoculum
3.1 Impact of a cost or of a benefit of resistance
In the main text, we considered the case where the resistant mutant is neutral in the absence of drug. However, antibiotic resistance often carries a cost in the absence of drug [Andersson99, Andersson10]. Besides, if antibiotic is already present at relatively low doses in the environment before the addition of drug, it can induce a benefit of resistance before drug is added. Here, we briefly discuss these two cases.
Cost of resistance.
A cost of resistance biases the competition between mutants and wildtypes, making it difficult for mutants to fix once they have appeared. We show in Figure S1 the survival probability of a spatially structured population after drug treatment, in the case where the fitness of the mutant is $f_{R}=0.9$ before drug addition (while $f_{S}=1$). We consider both a clique (i.e. the minimal model of spatial structure considered in most of the present work), and a star with different migration asymmetries $\alpha$. These results can be compared to those in the cost-free resistance case, shown in Figure 2 for the clique, and in Figure S5 for the star. We consider a similar range of drug addition time $t_{add}$ as in those figures. In Figure S1, we do not observe a statistically significant impact of population structure on the survival probability of the population.
With a cost of resistance $\delta=(f_{S}-f_{R})/f_{S}>0$ (giving $\delta=0.1$ in Figure S1) satisfying $\delta\ll 1$ and $N\delta\gg 1$ (satisfied in Figure S1), the probability of fixation of one mutant in a deme of size $N$ is approximately $\delta e^{-N\delta}$ in the diffusion approximation. It is thus exponentially suppressed, and the time to appearance of a successful mutant would be much longer than in the neutral case, on average of $\langle t_{af}\rangle\approx e^{N\delta}/(N\mu g\delta)$, giving $\langle t_{af}\rangle\approx 2.4\times 10^{13}$ with the parameter values of Figure S1. Therefore, fixation events can be neglected in the range of $t_{add}$ considered in Figure S1. Thus, we can estimate the survival probability from the probability of presence of mutant lineages destined for extinction, introduced in Equation (S16). This analytical prediction agrees with our simulation result, see Figure S1.
Still for resistant mutants with cost $\delta$ satisfying $\delta\ll 1$ and $N\delta\gg 1$, the average fixation time $\langle t_{af}\rangle\approx e^{N\delta}/(N\mu g\delta)$ of a successful mutant increases as $N$ increases. Thus, the mean fixation time of a successful mutant is substantially shorter in an average deme than in a well-mixed population. Therefore, if times comparable to the mean fixation time of a successful mutant in a deme (or larger) were considered, we predict that spatial structure would favor population survival even more than in the neutral case. While this is a potentially interesting effect, we do not consider it further here because of the extremely long times involved.
Figure S1: Survival probability of a structured bacterial population with a cost of resistance. The survival probability of the population to the addition of biostatic drug is shown versus the drug addition time. We consider a clique and a star with different values of migration asymmetry $\alpha$. The case of the well-mixed population with the same total size (“WMP”) is shown for reference. The migration rate is $\gamma=10^{-6}$ in the clique, while in the star we follow the convention in Section 4 to compare with a clique of a given migration rate $\gamma$, leading to $\gamma_{O}=(D-1)\gamma$ and $\gamma_{I}=\alpha\gamma_{O}$. Simulation results are obtained from $10^{3}$ replicates, with error bars representing 95% confidence intervals. Red horizontal dashed line (“Theory”): analytical prediction from Equation (S16). Parameter values: $K=200$, $D=5$, $f_{S}=1$ without drug, $f_{S}=0$ with drug, $f_{R}=0.9$ with and without drug, $g=0.1$, $\mu=10^{-5}$.
Benefit of resistance.
Let us now consider the case where the resistant mutant has a selective advantage $s=(f_{R}-f_{S})/f_{S}>0$ in the absence of drug. Assuming $Ns\gg 1$, the probability of fixation of one mutant in a deme of size $N$ is approximately $s$ in the diffusion approximation. As this is independent of $N$, the same holds for a well-mixed population. Thus, in this case, the average time to appearance of a successful mutant is the same in a deme and in a well-mixed population. This entails that the effect of spatial structure we evidenced for neutral and effectively neutral resistant mutants does not extend to beneficial ones.
3.2 Stochastic extinction
The probability of stochastic extinction starting from $j_{0}$ individuals in a population can be computed by formally integrating the Master equation with the initial condition $j=j_{0}$:
$P_{0}(t)=\left(e^{\mathbf{R}t}\right)_{0j_{0}},$ (S19)
where $\mathbf{R}$ is the transition rate matrix describing the population dynamics, see Ref. [marrec_resist_2020]. In particular, for a population with a carrying capacity $K=10^{2}$, the probability of rapid initial extinction starting from 10 individuals or more is negligible (lower than $10^{-9}$) [marrec_resist_2020]. This is why, in our categorization of results used in Figure 3, we called “Large R population” the case where the number of mutants is at least 10: the possibility of stochastic extinction can be neglected in this case.
3.3 Growth of mutant population in structured and well-mixed populations
Starting from an inoculum of sensitive bacteria, resistant mutants may appear and fix in the absence of drug. In the case of a structured population with small migration rates, they locally fix deme after deme. This results in a substantial difference in the growth pattern of the number of resistant individuals between a well-mixed and a structured population. We present one specific trajectory of the number of sensitive and resistant individuals in a well-mixed population in Figure S2A and one in a structured population in Figure S2B. In the structured population, a locally successful mutant appears on average earlier than a successful one in the well-mixed population. However, it then takes more time for resistance to fix in the whole population, as this involves migration of resistant mutants from deme to deme (or independent appearance of other mutants if migration rates are extremely small). Accordingly, Figure S2A features a rapid but late growth of mutant fraction in the well-mixed population. Meanwhile, Figure S2B, we observe that the growth of mutant fraction in the structured population features $D-1$ intermediate plateaus, as fixation occurs deme after deme.
Despite these differences in the individual trajectories of the number of mutants, their average over many replicates is the same whatever the migration rate, see Figure 3A. However, these differences in trajectories are at the root of those observed in the variance across replicates shown in Figure 3B. The well-mixed population features more variance in mutant numbers across replicates than structured populations with small migration rates, because of the variability of the appearance time of a successful mutant. More precisely, the appearance time of a successful mutant in the fastest deme of a structured population has smaller variance, as it is exponentially distributed with a larger rate, see Equation S9, compared to Equation S5 which holds for a well-mixed population.
Figure S2: Dynamics of population in a well-mixed and a structured population. We present one trajectory from a single simulation realization in the absence of drug for a well-mixed population in Panel A, and one for a structured population with same total size in Panel B. In both cases, we show the number of sensitive (S) and resistant (R) individuals versus time. Parameter values: $K=200$, $D=5$, $\mu=10^{-5}$, $f_{S}=f_{R}=1$, $g=0.1$, $\gamma=10^{-7}$.
3.4 Population composition as a function of time for different migration rates
In Figure 3C, we analyzed the composition of the population at the time featuring the largest inter-replicate variance for the well-mixed population, namely $t=5\times 10^{5}$, in spatially structured populations with different migration rates. In Figure S3, we report how this composition evolves in time, in the absence of drug, for spatially structured populations with different migration rates. As in Figure 3C, we focus on the number of mutants in each deme to categorize the population composition.
Figure S3: Population composition versus time for different population structures. Population composition in the absence of drug in the well-mixed population (“WMP”, Panel A), in the fully subdivided population (“SP”, Panel B), and in the clique-structured population with same total size for three values of the migration rate $\gamma$ (Panels C-E). The four categories of population composition reported here are based on the number of mutants per deme. They are defined in the main text and used in Figure 3C. The x-axis tick corresponding to $t=5\times 10^{5}$, considered in Figure 3C, is highlighted in purple. Data is obtained from $10^{4}$ simulation replicates in each case. Parameter values (in all panels): $K=100$, $D=10$, $f_{S}=f_{R}=1$, $g=0.1$, $\mu=10^{-5}$.
Recall that here, the population is composed of wild-type sensitive bacteria and of neutral resistant mutants that arise with probability $\mu$ upon division of wild-types. Figure S3 shows that when the per capita migration rate $\gamma$ is decreased, there is a longer phase where the population has an intermediate composition, with some demes harboring a substantial number of resistant mutants (“Big R pop.” category). This is consistent with our above discussion in Section 3.3 and with our results in Figure S2. The fixation of mutants occurs gradually, deme by deme, in structured populations with small migration rates. Furthermore, substantial mutant numbers exist earlier in these structured populations. We further observe in Figure S3 that a per capita migration rate $\gamma=10^{-3}$ results in an evolution of population composition akin to that of a well-mixed population, a per capita migration rate $\gamma=10^{-7}$ yields an evolution of population composition similar to that of a fully subdivided population with $\gamma=0$. Structures with intermediate values of $\gamma$ feature an intermediate evolution of population composition.
3.5 Dynamics at the single-deme level
In Figure 3, we analyzed the time evolution of mean and variance across replicates of the total number of mutants in the population. In Figure S4, we analyze these quantities at the single-deme level. Since the clique is a symmetric structure such that all demes are equivalent, we select deme 1 in each of our simulation replicates for this analysis. Figure S4A shows the average number of mutants in this deme as a function of time. This quantity is the same in the structures with different migration rates, as is the case at the population level (see Figure 3A). Conversely, Figure S4B shows that the variance of the number of mutants in one deme differs from its population-level counterpart (see Figure 3B). Indeed, at the deme level, there is no visible difference between different population structures as far as the variance across replicates of the number of mutants in a deme is concerned, see Figure S4B. Thus, the local mutant dynamics are the same for all migration rates. It is the global dynamics at the population level, combining all demes, which gives rise to the impact of spatial structure that we evidenced in this work.
Figure S4: Mean and variance of the number of mutants at the single-deme level. Panel A: the mean number of mutants in deme 1 is shown as a function of time in the absence of drug, for different values of the migration rate. Panel B: the variance across simulation replicates of the number of mutants in deme 1 is shown as a function of time in the absence of drug, for different values of the migration rate. Data is obtained from $10^{4}$ simulation replicates in each case. The case of an isolated deme, i.e. of a well mixed population with carrying capacity $K$ (“WMP ($K=100$)”), is shown for reference. Parameter values (in both panels): $K=100$, $D=10$, $f_{S}=f_{R}=1$, $g=0.1$, $\mu=10^{-5}$.
4 Lattice and star structures with sensitive inoculum
Our work mainly focuses on a minimal model of spatially structured populations where all demes are equivalent and connected to one another by identical migration rates. This structure is known as the island model, the clique or the fully connected graph. Let us now extend our study to different graph structures. We first consider the square lattice (see Figure S5A). It is a symmetric structure, like the clique, but migrations from each deme are restricted to its four nearest neighbors. More precisely, in the lattice, as in the clique, in each deme the inward migrations balance the outward migrations, i.e. for each $i$ we have $\sum_{j}m_{ij}=\sum_{j}m_{ji}$, where $m_{ij}$ denotes the migration rate from deme $i$ to deme $j$: this means that the lattice is a circulation [lieberman_evolutionary_2005, marrec_toward_2021, abbara_frequent_2023]. The per capita migration rate in the lattice is denoted by $\gamma_{L}$, equal in all directions. We further consider the star (see Figure S5B), which is less symmetric in the sense that the central deme is different from the leaf demes. Migrations from a leaf to the center occur at per capita rate $\gamma_{I}$, while migrations from the center to the leaf occur at per capita rate $\gamma_{O}$. We define the migration asymmetry parameter as $\alpha=\gamma_{I}/\gamma_{O}$. If $\alpha=1$, the star is a circulation, while for all other values of $\alpha$ it is not, which impacts the fixation probability of a mutant in the whole population [marrec_toward_2021, abbara_frequent_2023].
Figure S5: Survival probability of a bacterial population to biostatic drug, for different spatial structures. Panel A: Lattice structure with $D=9$ and migrations $\gamma_{L}$ to each nearest neighbor. Panel B: Star structure with $D=9$ and migration rates $\gamma_{O}$ from the center to a leaf and $\gamma_{I}=\alpha\gamma_{O}$ from a leaf to the center. Panel C: Survival probability of a population starting from a sensitive inoculum versus drug addition time $t_{add}$, for different spatial structures (fully subdivided population with no migrations, “$\gamma=0$”; lattice; clique; star with different migration asymmetries). Parameter values: $D=16$, $K=100$, $f_{R}=1$, $f_{S}=1$ before drug addition, $f_{S}=0$ when antibiotic is added, $g=0.1$, $\mu=10^{-5}$, $\gamma=10^{-6}$ for the clique, $\gamma_{L}=3.75\times 10^{-6}$ for the lattice, $\gamma_{O}=15\times 10^{-6}$ for the star, and $\gamma_{I}=\alpha\gamma_{O}$. Each result is obtained from $10^{3}$ simulation replicates. Error bars show the 95% confidence intervals.
Mutant fixation in the fastest deme is at the root of the impact of population structure on resistance evolution that we evidenced here, in particular in Figure 2. Suppose that a successful mutant appears in the fastest deme: wild-type individuals from connected demes can migrate to the fastest deme and take over there, which takes the population back to the initial fully wild-type state. Thus, to choose migration rates that make different structures comparable, we set migration rates so that the rate of invasion of one deme by the connected demes is the same for two structures. For the lattice, this leads to $\gamma_{L}=\gamma(D-1)/4$, where $\gamma$ is the per capita migration rate in the clique. Note that this condition also yields an equal total exchange rate in the clique and the lattice. For the star, we first remark that a successful mutant is $(D-1)$ times more likely to first appear in a leaf than in the center. Neglecting the case where a successful mutant first appears in the center leads to the condition $\gamma_{O}=(D-1)\gamma$. Note that this condition does not coincide with setting equal overall exchanges between demes in the star and in the clique.
Figure S5C shows the survival probability of bacterial populations with different spatial structures. We observe that the fully subdivided population with $\gamma=0$ has a larger survival probability than all others. This generalizes our results in Figure 2. It can be attributed to the fact that wild type bacteria cannot re-invade the fastest deme in the absence of migrations. Moreover, we find that the lattice and the clique with the same invasion rate do not feature any statistical difference in their survival probabilities. More generally, we observe that the specific graph structure of the population does not significantly affect the survival probability if the addition time $t_{add}$ of antibiotic satisfies $t_{add}\lesssim\langle t_{afF}\rangle$, where $\langle t_{afF}\rangle$ is the average time it takes for a successful mutant to appear in the fastest deme. The minor differences observed between the stars with different asymmetries and between them and other structures for $t_{add}>\langle t_{afF}\rangle$ can be attributed to the fact that more than one deme may then have fixed mutants. Recall indeed that our matching condition for re-invasion was constructed assuming that only one deme was mutant – and that for the star it did not match the condition of same total exchanges between demes.
5 Clique population structure with mutants in the inoculum
So far, we focused on a sensitive inoculum and on resistant individuals appearing through mutations. Let us now consider the case where mutants are already present in the inoculum. To isolate their effect, we do not include any new mutations in this case. This situation is highly relevant, especially because many experimental studies start from two bacterial strains, of which one is resistant to a given antibiotic and the other is not. In particular, a spatially structured population with mixed inoculum in each deme was recently studied [kreger_role_2023].
In Figure S6, we show the survival probability of a bacterial population to biostatic antibiotic treatment versus the addition time $t_{add}$ of antibiotic for different migration rates, in the case where each deme is inoculated with 5% mutants. Our simulation results show that the survival probability decreases when $t_{add}$ is increased in this scenario. This was expected, because new mutants do not appear, and the mutants initially present in the inoculum may go extinct. However, the spatial structure of the population still plays a crucial role. Indeed, as the migration rate $\gamma$ decreases, the survival probability increases. This is reminiscent of our result with a sensitive inoculum (see Figure 2). We added the expected probabilities to still have mutants in the system as dashed lines. In a well-mixed population with initial fraction of mutants $x_{ini}$, the probability of having mutants in the long term is equal to their fixation probability, i.e. to $x_{ini}$ since we are considering neutral mutants. In a fully subdivided population with no migrations, this probability is equal to $1-(1-x_{ini})^{D}$, which is the probability that there is at least one deme where mutants fix. In a spatially structured population with nonzero migration rates, the survival probability gradually converges to the well-mixed population one because of migrations.
Figure S6: Survival probability of a structured bacterial population with mutants in the inoculum upon addition of a biostatic drug. The survival probability is plotted versus the addition time $t_{add}$ of antimicrobial. We consider a clique population structure with various migration rates. Results for a well-mixed population with same total size (“WMP”) are presented as reference. Horizontal dashed lines: analytical predictions of the probabilities to still have mutants in the system in the long term (dark blue: well-mixed population; yellow: fully subdivided population with $\gamma=0$). Data is obtained from $10^{4}$ replicate simulations in each case. Parameter values: $K=100$, $D=10$, $f_{S}=1$ without drug, $f_{S}=0$ with drug, $f_{R}=1$, $g=0.1$, $\mu=0$. Initial percentage of mutants in each deme (and in the well-mixed population): 5%.
[…]
[heading=bibintoc]